OpenStudy (anonymous):

A Closed box with a square base must have a volume of 10 cubic feet. Express the amount A of material used to make such a box as a function of the length x of a side of the square base.

OpenStudy (anonymous):

and then i bet we are going to find the least amount of material. put x = length of bottom edge so area of base is \[x^2\] since it is a square and area of top is also \[x^2\] since the box is closed. if the height of the box is say "h" then the volume of the box is \[x^2h\] but since you know that the box contains 10 cubic feet, this tells you that \[x^2h=10\] so \[h=\frac{10}{x^2}\]

OpenStudy (anonymous):

we have two squares both with side x to that takes \[2x^2\] of material. then we have 4 sides with area \[xh\] so that takes \[4xh\] material for a total of \[2x^2+4xh\] material. now replace h by \[\frac{10}{x^2}\] and we get \[A(x)=2x^2+\frac{40}{x}\] if my algebra is correct